OpenStudy (gabbyalicorn):

What is the surface area of a rectangular prism with a length of 6 in., a width of 4 in., and a height of 10 in.?

OpenStudy (michele_laino):

here we have to do this computation: area A is: A=2*(6*4+6*10+4*10)=...

OpenStudy (gabbyalicorn):

6080

OpenStudy (michele_laino):

here is another step: A=2*(24+60+40)=...

OpenStudy (gabbyalicorn):

248

OpenStudy (michele_laino):

that's right!

OpenStudy (gabbyalicorn):

:D yay! Thank you!

OpenStudy (michele_laino):

:)